HUBBLE SPACE TELESCOPE DAILY REPORT #2507 PERIOD COVERED: 0000Z (UTC) 11/23/99 - 0000Z (UTC) 11/24/99 Daily Status Report as of 328/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: Since HST remains in Zero Gyro Sun Point (ZGSP), no science observations were performed. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No activity 2.2 FHST Updates: No activity 2.3 Operations Notes: The HST continues in ZGSP. The SSR EDAC error counter was cleared once using ROP SR-1. DSN sites were again utilized to dump engineering data to the ground. HSTAR 7276 was written when NICMOS safed itself at 326/193618Z, shortly after the instrument had been recovered to boot mode. At 327/130802Z, side 2 of the STIS low voltage power supply S/C voltage flagged out-of-limits high at 33.26 volts. HSTAR 7375 was written. The Time Base Recharge Ratio Buffer was re-set at 327/1530Z. The NSSC-1 status buffer was dumped and re-set at 327/2039Z as directed by ROP NS-3. As documented in HSTAR 7377, at 327/144720Z, the NICMOS 28 volt bus voltage for side 1 (NC28VIN1) flagged our-of-limits high at 31.608030 volts. Per an operations note, this parameter was adjusted at 328/0334Z. Note: The next Daily Report will be issued on November 29, 1999. Enjoy the holiday! 3.0 SIGNIFICANT FORTHCOMING EVENTS: Servicing Mission 3A, currently scheduled for December 9, 1999. /CAW